OpenStudy (anonymous):

what is the derivative of \[\frac{ xsinx }{ x ^{2}+1 }\] in it's most simplified form

OpenStudy (anonymous):

please help i got a really long answer and I need to find the second derivative as well so I figured there must be a simplified form but I'm struggling to get it

OpenStudy (anonymous):

\[f(x)=\frac{x\sin x}{x^2+1}\\ \begin{align*}f'(x)&=\frac{(x^2+1)(\sin x+x\cos x)-(x\sin x)(2x)}{(x^2+1)^2}\\ &=\frac{(x^3+x)\cos x+(1-x^2)\sin x }{(x^2+1)^2}\end{align*}\]

OpenStudy (anonymous):

thanks a bunch